I have worked out the answer to the question but cannot explain the logic to my 9 year old son. His teacher has explained how to do it. Can you please help. The sum of two numbers is 147. The difference between the two numbers is 43. What are the two numbers.
Answer by nadinadan(32)

So, we need to change the words into mathematical operations.The problems says:
A+B=147 and A-B=43, so we take the second equation and write it down:
A-B=43 so we need to find what A is A=43+B
We replace the A from this ecuation into the first ecuation so instead of A+B=147, we will have 43+B+B=147 so 2B=147-43 and 2B=104 so B=104/2 which equals 52. Now we replace B in any of these ecuations with the value 52 and solve the ecuation to find A. The rule is from one ecuation find one of the unknown values and replace it into the other ecuation.
Try this way, it may work.